An antacid tablet containing 1.25 g NaHCO3 is dissolved in water and the total volume of solution is
500.0 mL. Calculate the molarity of this NaHCO3 solution.
1
Expert's answer
2021-09-12T00:43:15-0400

CM = n / V

n = m / M

M (NaHCO3) = 84.01 g/mol

n (NaHCO3) = 1.25 / 84.01 = 0.015 mol

CM (NaHCO3) = 0.015 / 0.5000 = 0.03 M